# Towards a Deeper Understanding of Reasoning Capabilities in Large Language Models

### v1

@@ 5 Conclusion and Discussion
Our evaluation of open-source LLMs on the SmartPlay benchmark reveals both the potential and limitations of advanced prompting strategies, including self-reflection, heuristic mutation, and planning We find that excessive reasoning harms the performance of smaller models on simple tasks, as it forces the model to filter through more content and increase the signal-to-noise-ratio [26, 11, 9]. This reasoning not only distracts [26, 11] but also causes the models to overthink [2, 24, 5], leading the model to overcomplicate the process and disregard simpler, more effective solutions. In addition, we find that larger models perform better, but that strategic prompting can close this gap, and that having a dense, task-aligned reward signal can improve an agent’s decision-making, offering a simpler alternative compared to the significant effort needed to find optimal prompts. While smaller models particularly benefit from advanced prompting in complex tasks, our results demonstrate significant variability: the same prompt can produce substantial gains or, conversely, lead to worse performance than the baseline. This highlights the brittleness of current techniques and the need for more robust solutions. Moreover, the common practice in reasoning studies of reporting aggregate performance metrics, such as accuracy [30, 1, 10] or F1 scores [34] without including variability measures can be misleading and obscure the sensitivity of results to prompt variations caused by randomness or instability, affecting the generalization and reproducibility of findings.
Finally, while LLMs might exhibit proficiency on in-distribution data, we find little evidence for emergent reasoning or self-learning. Frequent failure modes include hallucinating invalid action trajectories and getting stuck in loops. Our findings support the growing evidence necessitating a critical examination of prompting methods claiming emergent LLM abilities [28] and highlight the need to re-evaluate current benchmarks like question-answer pairs or math word problems, which inadequately capture the complexity of reasoning and fail to reveal inherent deficits [15]. Future work could benefit from combining in-context learning with external memory to improve recall, symbolic abstractions to ensure verifiable reasoning, and multimodal perception to ground the agent’s understanding in the physical world.

### v2

@@ V Conclusion and Discussion
Our evaluation of open-source LLMs on the SmartPlay benchmark reveals both the potential and limitations of advanced prompting strategies, including self-reflection, heuristic mutation, and planning. We find that excessive reasoning harms the performance of smaller models on simple tasks, as it forces the model to filter through more content and increase the signal-to-noise ratio [17]. This reasoning not only distracts [17] but also causes the models to overthink [19], leading the model to overcomplicate the process and disregard simpler, more effective solutions. In addition, we find that larger models perform better, but that strategic prompting can close this gap, and that having a dense, task-aligned reward signal can improve an agent’s decision-making, offering a simpler alternative compared to the significant effort needed to find optimal prompts. While smaller models particularly benefit from advanced prompting in complex tasks, our results demonstrate significant variability: the same prompt can produce substantial gains or, conversely, lead to worse performance than the baseline; this holds for all model sizes. This highlights the need for more robust solutions. Moreover, the common practice in reasoning studies of reporting aggregate performance metrics, such as accuracy [9] or F1 scores [22] can be misleading and obscure the sensitivity of results to variations caused by instability, affecting the generalization and reproducibility of findings. Finally, while LLMs might exhibit proficiency on in-distribution data, we find little evidence for emergent reasoning. We find that LLMs expose a knowing-doing gap[23]: models can recite optimal solution learning during training (e.g. the seven-move Tower of Hanoi) but fail to execute them in practice. We also find a language-embodiment gap: textual understanding does not translate into grounded, stateful planning. Our findings support the growing evidence necessitating a critical examination of prompting methods claiming emergent LLM abilities and highlight the need to re-evaluate current benchmarks like question-answer pairs or math word problems, which inadequately capture the complexity of reasoning and fail to reveal inherent deficits. Future work could benefit from combining in-context learning with external memory to improve recall, symbolic abstractions to ensure verifiable reasoning, denser task‑aligned feedback, and multimodal perception to ground the agent’s understanding in the physical world.